Mr. Bookman, a new fiction author, wishes to know whether there is any relationship between a book’s length and the price charged for the book. He chooses a random sample of 10 novels and records the length and price of each as follows:

Length (pages) 520 680 740 200 400 800 750 500 300 350
Price (Ksh 00) 16 20 21 9 10 22 20 15 10 10

a) Is there a significant relationship between novel length and its price? (use test for significance of correlation)

b) Develop a regression model for predicting the price of a novel using its length.

c) Test the significance of the model using t-test, R2 and F-test.
1
Expert's answer
2019-11-11T09:51:25-0500

a) the correlation coefficient is calculate using correlation function in Excel

"CORREL (array1, array2) = 0.977"

The following needs to be tested:

"Null: \rho = 0"

"Alternative:\rho\ne0"

The sample size = 10. This means DF = n-2 =10 -2 =8


The corresponding critical correlation value "r_c" for a significance level of  "\alpha =0.05", for a two-tailed test is:

"r_c =0.632"

Observe that in this case, the null hypothesis "H_0: \rho = 0" is rejected if "|r| >r_c =0.632"


we have that "|r| =0.977 > r_c=0.632" , from which is concluded that the null hypothesis is rejected. There is a significant correlation.


b) Regression from Excel toolPak has following output predicting prices





The regression model equation: "Price (Ksh) =2.424 length (pages) + 259.729"


c) Testing the model

"R^2" indicates there is goodness-of-fit because 95.44% of the variation on prices is contributed by the model. Overall, the model is statistically is significant since "F(1,8) =167.08, p-value < 0.05" . The regression coefficient for length is also statistically significant because the corresponding t-stat has a "p-value < 0.05."